# zkMixer: A Configurable Zero-Knowledge Mixer with Proof of Innocence and Anti-Money Laundering Consensus Protocols

### v1

@@ 3.2 System Components
3.2.1 Arithmetic Circuits
For the ZK functionality of our protocol we use Circom, [^3] https://iden3.io/circom a programming language used to build arithmetic circuits that can be used in ZKPs; and snarkjs, [^4] https://github.com/iden3/snarkjs a library implementing zkSNARK schemes. The arithmetic circuits are first compiled using Circom, and then snarkjs is used to create a smart contract that can verify the ZKPs for withdrawals on-chain. For our prototype, the Groth16 scheme was used as it is the most popular, but this can be easily substituted by other schemes supported by snarkjs, such as PlonK.
We have implemented three arithmetic circuits. The first circuit creates the commitment from the secret and the nullifier by applying the Poseidon Hash. The second circuit uses the first and is responsible for calculating the root of the Merkle tree; given a commitment, and the rest of the tree nodes in the path from the commitment to the root. Finally, the last circuit, which uses the previous two, is responsible for the ZKPs that are used to withdraw from the mixer. This circuit takes the user’s input, calculates the user’s commitment, calculates the root of the Merkle tree, and finally verifies that the root is the same as the root in the smart contract.
3.2.2 zkMixer Smart Contract
The zkMixer smart contract contains the core functionality of our mixer. It follows the design presented in Section 2.5, with some additional functionality. zkMixer contains a list of pre-deposits, i,e., deposits before they are added to the mixer’s Merkle tree, along with functionality to freeze and unfreeze them, to add them to the mixer, and to delete them and send the funds to a different address (either refund or confiscate them). The properties of the mixer, such as the mixer mode and the deposit amount, must be configured prior to deployment. As shown in Fig. 4, the mixer operates in one of four distinct modes, based on how pre-deposits are approved. The modes are as follows:
@@ 3.2.1 Arithmetic Circuits
3.2.3 Variable Lock Time
@@
3.2.4 multiSig Smart Contract
The multiSig smart contract contains all the functionality used by the verifiers. The smart contract has an owner who can add and remove verifiers. This can be a centralized authority, a DAO, or something similar depending on the intended use case. Verifiers are responsible for checking that the deposits conform to the rules of the mixer. These verifiers can be users of the mixer that want to participate in its governance or blockchain analytics companies that provide real-time screening services. [^5] Examples include Elliptic (https://www.elliptic.co/solutions/screening), TRM (https://www.trmlabs.com/blockchain-intelligence-platform/wallet-screening), Chainalysis (https://www.chainalysis.com/solution/crypto-compliance-assess/), and AnChain.AI (https://www.anchain.ai/ciso) The services of the verifiers can be rewarded by adding a fee to all deposits, although for simplicity this has not been implemented in our prototype. More advanced incentive and punishment mechanisms can also be implemented, but again for simplicity these were omitted from our prototype.
Any single verifier using the multiSig contract can freeze any of the pre-deposits to the mixer. Freezing a deposit indicates that they have potentially identified a deposit that does not conform to the requirements of the mixer. At that point, that deposit cannot be added to the mixer even if its lock time has elapsed. A verifier can then propose what should be done with the deposit and indicate an address to which the deposit should be sent. If the multiSig smart contract was configured to allow for confiscation of funds, the address can be any address, but if it was not, the address is set to the address of the user who deposited the funds, i.e., only refunds are allowed. The rest of the verifiers then vote on the proposal. If the majority vote for the proposal, then the pre-deposit is deleted from the mixer, and deposited funds are sent to the specified address. However, if the majority of the verifiers do not agree with the proposal, then the proposal fails, and the deposit is admitted to the mixer. MultiSig approvals work similarly.
@@
3.3 Use cases
@@
Figs. 5 and 6 show two examples of successful deposits. In Fig. 5, the user first makes a deposit to the mixer, which creates a pre-deposit on zkMixer. Then, after the lock time has passed, as the pre-deposit was not flagged, the user is allowed to call the approval function, which deletes the pre-deposit and adds the users’ commitment to the mixer’s Merkle tree. Some time later, the funds are withdrawn anonymously from the mixer by a different user. Fig. 6 shows a different example of successful use of the mixer. The user again makes a deposit, but in this case the verifiers use multiSig to approve the deposit by consensus, before a withdrawal takes place.
@@
Examples of failure cases are shown in Figs. 7 and 8. The former shows an example of a refund, while the latter shows an example of a confiscation. In both cases, the user makes a deposit to the mixer, which is found suspicious by a verifier and frozen using the multiSig. In the first example (Fig. 7), the verifier makes a proposal to reject the deposit and refund the user. The rest of the verifiers reach consensus and approve the proposal. The funds are then transferred to the original sender, and the deposit is removed from the smart contract. In the second example (Fig. 8), the verifier proposes a confiscation, which is approved by consensus. In this case, the funds are sent to a different user, who is known to be the rightful owner of the stolen funds.
4 Analysis and Discussion
4.1 Costs of Deployment and Operations
@@
TABLE I: Gas and Monetary Cost of Contract Deployment
Smart Contract Gas Cost USD Cost*
@@
  • * If the system were to be used on the Ethereum mainnet, using a gas price of 1.00 Gwei and ETH price of \$2000.
TABLE II: Gas and Monetary Cost of Operations
Smart Contract Operation Gas Cost USD Cost*
@@
  • * If the system were to be used on the Ethereum mainnet, using a gas price of 1.00 Gwei and ETH price of \$2000.
4.2 Lock Time
The lock time is one of the most important properties of our protocol. Many variables, such as the capabilities of the validator and the expected traffic to the mixer, must be considered before an appropriate lock time is selected. In general, for safety, higher lock times should be preferred as they allow more time for verification. If lock times are too short, it can lead to verifiers being overwhelmed and incapable of adequately scrutinizing deposits. This opens a potential exploit for malicious actors, where large amounts of deposits are deliberately submitted in quick succession, with the aim of causing validators to have insufficient time to vet all deposits before the lock time runs out.
@@ 4.2 Lock Time
4.3 Network of Verifiers
@@
The current prototype uses a consensus mechanism implemented by a multiSig smart contract. This is simple and provides the desirable functionality, but might not be ideal for a real-world deployment. An alternative solution, such as Decentralized Oracle Networks (DONs), implemented by Chainlink [50], could be used instead. Chainlink and other similar solutions provide a decentralized network of oracles that can act as the validators in our protocol. The oracles in DONs have the necessary knowledge and experience to run mission-critical infrastructure, and are trusted to act as the backbone of many DeFi protocols.
4.4 Deposit Verification
@@
We assume that the network of verifiers will be able to identify fraudulent deposits to the mixer. This is a difficult task, and if the verifiers are not up to it, then the whole protocol fails. We believe that verification can be achieved using a combination of self-reporting from exploited parties and dedicated nodes running tracing programs. When a hack is noticed, the affected party can report it to a registry. Then the trackers will be able to maintain a banned list of all the addresses that the funds move through. If any of those addresses attempts to interact with the mixer, it can be flagged by the verifiers.
4.5 Interactions between Mixers
@@
The protocol we have presented can be configured to suit the needs of different users and businesses. We imagine that this will lead to many instances of mixers operating under different rule sets. For example, a mixer operating in the US might base its rules on the OFAC list of sanctioned addresses, while one operating in Russia might use a different list. This creates a problem as different rule sets might be incompatible, resulting in a situation where different mixers may reject outputs from each other; e.g., a user who has previously received funds from the Russian mixer might be unable to deposit into the US mixer.
@@
We recognize this problem, but we believe that our protocol provides a solution. In our example, a third mixer (which we will call “High Bar”) can be set up to require verification of all deposits, and there will be very high requirements on deposits before they are allowed. The High Bar mixer will be operated by verifiers who can be contacted by users off-chain and who are willing to spend time auditing the source of the user’s funds (such verifiers could be a government department or national regulator, for instance). If the audit shows that the funds received from the Russian mixer are legitimate and do not originate from illicit activities, then the user will be able to pass these funds through the High Bar mixer. Due to the increased level of scrutiny provided by High Bar, all other US-based mixers will accept its outputs as legitimate funds.
4.6 Fracturing of Anonymity Pools
@@
One of the benefits of existing mixers is the large anonymity sets. In existing mixers, the pools of funds are usually only broken to allow for different deposit amounts. For example, Tornado Cash had four pools corresponding to 0.1, 1, 10, and 100 Eth. In our protocol, the pools of funds are not only broken to allow for different amounts; pools are also based on user requirements, legislation, use cases, etc. This comes at the expense of fracturing the anonymity pools of users, thus reducing the privacy protection of the system.
@@
5 Conclusion
In this paper, we examined cryptocurrency mixers, we categorized them based on their characteristics, we detailed how they operate, and we presented evidence of their use in money laundering. We then discussed Proof of Innocence (PoI), a solution previously proposed to prevent money laundering in mixers, and identified why PoI cannot work in practice. To address these issues, we introduced a novel mixer protocol that mitigates the PoI problem by using a consensus mechanism. Our protocol design allows groups of users to set up different instances of the mixer, each governed by the requirements that the group wants to achieve. Analysis of a system prototype shows that operational and deployment costs are low, with deployment costing around \$20 and each operation costing less than \$1. We release a full prototype of our protocol, which we have made open-source available at https://github.com/lifeisbeer/zkMixer.
Acknowledgments

### v2

@@ III-C System Components
III-C1 Arithmetic Circuits
For the ZK functionality of our protocol we use Circom, [^4] https://iden3.io/circom a programming language used to build arithmetic circuits that can be used in ZKPs; and snarkjs, [^5] https://github.com/iden3/snarkjs a library implementing zkSNARK schemes. The arithmetic circuits are first compiled using Circom, and then snarkjs is used to create a smart contract that can verify the ZKPs for withdrawals on-chain. For our prototype, the Groth16 scheme was used as it is the most popular, but this can be easily substituted by other schemes, such as PlonK.
Using Groth16 requires a circuit-specific trusted setup and a one-time procedure to generate public parameters, which also generates an unwanted trapdoor as a byproduct [43]. This trapdoor can be used to create fraudulent proofs; therefore users must trust that this procedure is done in a way that no single party has control over the trapdoor or that the trapdoor was discarded by the creator of the system. Groth16 could be substituted with a different proving system (such as PlonK) that only requires a universal trusted setup. Several universal trusted setup ceremonies have been performed by others using secure multiparty computation (e.g., Semaphore), [^6] https://github.com/privacy-scaling-explorations/perpetualpowersoftau which eliminates trust in the creator of this system.
We have implemented three arithmetic circuits. The first circuit creates the commitment from the secret and the nullifier by applying the Poseidon Hash. The second circuit, uses the first and is responsible for calculating the root of the Merkle tree given a commitment, and the rest of the tree nodes in the path from the commitment to the root. Finally, the last circuit, which uses the previous two, is responsible for the ZKPs that are used to withdraw from the mixer. This circuit takes the user’s input, calculates the user’s commitment, calculates the root of the Merkle tree, and finally verifies that the root is the same as the root in the smart contract.
III-C2 zkMixer Smart Contract
The zkMixer smart contract contains the core functionality of our mixer. It follows the design presented in Section II-D, with some additional functionality as described in Section III-B. zkMixer contains a list of pre-deposits, i.e., deposits before they are added to the mixer’s Merkle tree, along with functionality to freeze and unfreeze them, to add them to the mixer, and to delete them and send the funds to a different address (either refund or confiscate them). The properties of the mixer, such as the mixer mode and the deposit amount, must be configured prior to deployment. As shown in Fig. 4, the mixer operates in one of four distinct modes, based on how pre-deposits are approved. The modes are as follows:
@@ III-C1 Arithmetic Circuits
III-C3 Variable Lock Time
@@
III-C4 multiSig Smart Contract
The multiSig smart contract contains all the functionality used by the verifiers. The smart contract has a manager who can add and remove verifiers. This can be a centralized authority, a DAO, or something similar depending on the intended use case. Verifiers are responsible for checking that the deposits conform to the rules of the mixer. These verifiers can be users of the mixer that want to participate in its governance or blockchain analytics companies that provide real-time screening services. [^7] Examples include Elliptic (https://www.elliptic.co/solutions/screening), TRM (https://www.trmlabs.com/blockchain-intelligence-platform/wallet-screening), and Chainalysis (https://www.chainalysis.com/solution/crypto-compliance-assess/) The services of the verifiers can be rewarded by adding a fee to all deposits, although for simplicity this has not been implemented in our prototype. More advanced incentive and punishment mechanisms can also be implemented, but again for simplicity these were omitted from our prototype.
@@
Any single verifier using the multiSig contract can freeze any of the pre-deposits to the mixer. Freezing a deposit indicates that they have potentially identified a deposit that does not conform to the requirements of the mixer. At that point, the deposit cannot be added to the mixer even if its lock time has elapsed. A verifier can then propose what should be done with the deposit and indicate an address to which the deposit should be sent. If the multiSig smart contract was configured to allow for confiscation of funds, the address can be any address, but if it was not, the address is set to the address of the user who deposited the funds, i.e., only refunds are allowed. The rest of the verifiers then vote on the proposal. If the majority vote for the proposal, then the pre-deposit is deleted from the mixer, and deposited funds are sent to the specified address. However, if the majority of the verifiers do not agree with the proposal, then the proposal fails and the deposit is admitted to the mixer. MultiSig approvals work similarly.
III-D Use cases
@@
Fig. 5 and Fig. 6 show two examples of successful deposits. In Fig. 5, the user first makes a deposit to the mixer, which creates a pre-deposit on zkMixer. Then, after the lock time has passed, as the pre-deposit was not flagged, the user is allowed to call the approval function, which deletes the pre-deposit and adds the user’s commitment to the mixer’s Merkle tree. Some time later, the funds are withdrawn anonymously from the mixer by a different user. Fig. 6 shows a different example of successful use of the mixer. The user again makes a deposit, but in this case the verifiers use multiSig to approve the deposit by consensus, before a withdrawal takes place.
@@
Examples of failure cases are shown in Fig. 7 and Fig. 8. The former shows an example of a refund, while the latter shows an example of a confiscation. In both cases, the user makes a deposit to the mixer, which is found suspicious by a verifier and frozen using the multiSig. In the first example (Fig. 7), the verifier makes a proposal to reject the deposit and refund the user. The rest of the verifiers reach consensus and approve the proposal. The funds are then transferred to the original sender, and the deposit is removed from the smart contract. In the second example (Fig. 8), the verifier proposes a confiscation, which is approved by consensus. In this case, the funds are sent to a different user, who is known to be the rightful owner of the stolen funds.
IV Analysis and Discussion
IV-A Costs of Deployment and Operations
@@
TABLE I: Gas (Gwei) and Monetary (USD) Cost of Contract Deployment
Smart Contract Gas Cost \$ Cost
@@
TABLE II: Gas (Gwei) and Monetary (USD) Cost of Operations
Smart Contract Operation Gas Cost \$ Cost
@@
IV-B Lock Time
The lock time is one of the most important properties of our protocol. Many variables, such as the capabilities of the verifiers and the expected traffic to the mixer, must be considered before an appropriate lock time is selected. In general, for safety, higher lock times should be preferred as they allow more time for verification. If lock times are too short, it can lead to verifiers being overwhelmed and incapable of adequately scrutinizing deposits. This opens a potential exploit for malicious actors, where large amounts of deposits are deliberately submitted in quick succession, with the aim of causing verifiers to have insufficient time to vet all deposits before the lock time runs out.
@@ IV-B Lock Time
Longer lock times could potentially make the system less user-friendly. That being said, in order to prevent deanonymization, mixer users should ideally retain their deposits in the mixer for some time before withdrawing them. Tornado Cash, in a blog post, recommended that its users wait for at least five additional deposits after their deposit, and also for at least 24 hours, before withdrawing [44]. The rationale is that if there is a deposit followed immediately by a withdrawal, then an observer could guess that the two originate from the same user. Therefore, introducing a lock time should not negatively affect the usability of the mixer.
IV-C Network of Verifiers
@@
The current prototype uses a consensus mechanism implemented by a multiSig smart contract. This is simple and provides the desirable functionality, but might not be ideal for a real-world deployment. Such a system requires incentives for participants and thus could be very expensive in practice. In addition, it introduces a trust source and is susceptible to collusion. The manager and verifiers have too much control over the system and could act maliciously, either denying user access or even stealing funds. An alternative solution, such as Decentralized Oracle Networks (DONs), could be used instead. The oracles in DONs have the necessary knowledge and experience to run mission-critical infrastructure, and are trusted to act as the backbone of many DeFi protocols.
In any case, the verifiers could be existing transaction screening service providers. These providers already identify and track fraudulent activity on chain and so could easily monitor the deposits to zkMixer. Real-time AI-based scoring systems could also be used to provide a risk score for each deposit [36].
IV-D Deposit Verification
@@
We assume that the network of verifiers will be able to identify fraudulent deposits to the mixer. This is a difficult task, and if the verifiers are not up to it, then the whole protocol fails. We believe that verification can be achieved using a combination of self-reporting from exploited parties and dedicated nodes running tracing programs. When a hack is noticed, the affected party can report it to a registry. An example of such reporting is done by Bybit for the recent hack they suffered. [^8] https://www.lazarusbounty.com/en/ Then the nodes will be able to maintain a banned list of all the addresses that the funds move through. If any of those addresses attempts to interact with the mixer, it can be flagged by the verifiers.
IV-E Interactions between Mixers
@@
This protocol can be configured to suit the needs of different users and businesses. We imagine that this will lead to many instances of mixers operating under different rule sets. For example, a mixer operating in the US might base its rules on the OFAC list of sanctioned addresses, while one operating in Russia might use a different list. This creates a problem as different rule sets might be incompatible, resulting in a situation where different mixers may reject outputs from each other; e.g., a user who has previously received funds from the Russian mixer might be unable to deposit into the US mixer.
@@
We recognize this problem, but we believe that our protocol provides a solution. In our example, a third mixer (which we will call “High Bar”) can be set up to require verification of all deposits, and there will be very high requirements on deposits before they are allowed. The High Bar mixer will be operated by verifiers who can be contacted by users off-chain and who are willing to spend time auditing the source of the user’s funds (such verifiers could be a government department or national regulator, for instance). [^9] Similar to applications to OFAC for licenses to otherwise prohibited transactions (https://ofac.treasury.gov/ofac-license-application-page) If the audit shows that the funds received from the Russian mixer are legitimate and do not originate from illicit activities, then the user will be able to pass these funds through the High Bar mixer. Due to the increased level of scrutiny provided by High Bar, all other US-based mixers will accept its output as legitimate funds.
IV-F Fracturing of Anonymity Pools
@@
One of the benefits of existing mixers is the large anonymity sets. In existing mixers, the pools of funds are usually only broken to allow for different deposit amounts. For example, Tornado Cash had four pools corresponding to 0.1, 1, 10, and 100 Eth. In our protocol, the pools of funds are not only broken to allow for different amounts; pools are also based on user requirements, legislation, use cases, etc. This comes at the expense of fracturing the anonymity set of users, thus reducing the privacy protection of the system.
@@
IV-G Concrete Instantiation
For completeness, we include a practical, concrete example of how the protocol could be deployed. Lock time: we believe that a variable lock time is best for managing periods of high traffic. Non-confiscation: we encourage against allowing verifiers to confiscate deposits. Confiscation not only opens the protocol to an additional attack vector, but also limits its reach, as many users will avoid it due to the lack of custody during deposits. Verifiers: we suggest that the system has a small number of verifiers (e.g., 3), each of which should be an established screening company. This not only reduces fees, but, as the verifiers cannot confiscate deposits, it does not severely impact security. Funds tracking: finally, we would expect such a protocol to only track funds from relatively large hacks or thefts, as it is much more difficult for verifiers to adjudicate on small instances of theft, such as a dispute between two users. This, along with self-reporting, can make tracking feasible.
In Table III, we compare our concrete instantiation against a ZK mixer using PoI. As can be seen, the concrete instantiation is slower, has higher fees, and is susceptible to denial-of-service (DoS) attacks if the validators are malicious. However, the concrete instantiation provides better compliance as it avoids issues with PoI, and instead uses transaction screening to prevent bad actors from using the system. As explained in Section IV-B, speed is not critical for a mixer. Using only a few verifiers should also help minimize the additional cost. Therefore, DoS is the only major drawback of this concrete instantiation; but as all interactions are public on chain, it would be easy for users to prove that DoS is taking place. Overall, we believe that this is a worthwhile drawback given the compliance benefits.
TABLE III: Comparison between concrete instantiation (described in Section IV-G) and a ZK mixer using PoI
Properties Protocol Concrete Instantiation ZK Mixer using PoI
Speed Initial lock time Instantaneous
Cost Gas and validator fees Only gas fees
Trust Denial of serivce possible Zero trust
Compliance Deposit screening PoI can be circumvented
V Conclusion
In this paper, we examined cryptocurrency mixers, we categorized them based on their characteristics, we detailed how they operate, and we presented evidence of their use in money laundering. We then discussed Proof of Innocence (PoI), a solution previously proposed to prevent money laundering in mixers, and identified why PoI cannot work in practice. To address these issues, we introduced a novel mixer protocol that mitigates the PoI problem by using a consensus mechanism. Our protocol design allows groups of users to set up different instances of the mixer, each governed by the requirements that the group wants to achieve. We release a full prototype of our protocol, which we have made open-source available at https://github.com/lifeisbeer/zkMixer.
Acknowledgment